# Breaking Annotation Barriers: Generalized Video Quality Assessment via Ranking-based Self-Supervision

### v2

@@ 3.2.1 Pseudo-labeling using VQA Models
For soft ranking, we further compute the score variance for the video pair $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. Assuming the quality difference $\Delta=\overline{j}^{A}-\overline{j}^{B}$ follows a Gaussian distribution $\mathcal{N}(\Delta;0,\sigma^{2}_{\Delta})$, where $\sigma_{\Delta}=\sqrt{\sigma^{2}_{A}+\sigma^{2}_{B}}$, we assign labels based on statistical significance thresholds adapted from [77]: pairs are labeled as “superior” if $\Delta>2\sigma_{\Delta}$, “better” if $\sigma_{\Delta}<\Delta\leq 2\sigma_{\Delta}$, “similar” if $-\sigma_{\Delta}<\Delta\leq\sigma_{\Delta}$, “worse” if $-2\sigma_{\Delta}<\Delta\leq-\sigma_{\Delta}$, and “inferior” if $\Delta\leq-2\sigma_{\Delta}$.
3.2.2 Quality Ranking via Distortion Simulations
Figure 3: Overall framework of the proposed LMM-PVQA model. Compared to the standard LMM, our visual feature extractor adopts a dual-branch design with an additional motion encoder and projector to capture temporal-related distortions. The model takes a video pair and a text prompt as inputs and generates a text output specifying which video has better quality.
We introduce synthetic distortions to simulate typical degradation that may occur in real-world scenarios, which are categorized into three types: spatial distortions, temporal distortions, and streaming distortions. Specifically, spatial distortions include resolution resizing, Gaussian blur, Gaussian noise, darkening, and brightening, which simulate capture-related artifacts. Temporal distortions consist of jitter and stuttering, mimicking playback issues commonly observed in practical settings. Streaming distortions involve H.264 and H.265 compression, reflecting compression artifacts introduced by streaming media platforms. We denote this subset of data as PLVD-DS. The detailed generation of synthetic distortions are provided in the supplement file.
@@ 3.2.2 Quality Ranking via Distortion Simulations
We leverage distortion severity levels (e.g., constant rate factor for compression) as pseudo-labels to infer relative quality. Given a primary video $x^{0}$ and a synthetic distortion simulator $\mathcal{S}$, we degrade $x^{0}$ across $N_{\mathcal{S}}$ severity levels to generate distorted videos $\{x_{\mathcal{S}}^{i}\}^{N_{\mathcal{S}}}_{i=1}$. Pairs $(x_{\mathcal{S}}^{i},x_{\mathcal{S}}^{j})$ are randomly sampled. For hard ranking, pairs are directly annotated as “better” if $i<j$ or “worse” otherwise. For soft ranking, pairs with a severity difference $|i-j|>1$ are labeled as “superior” or “inferior” depending on the relative order of $i$ and $j$, while pairs with $|i-j|=1$ receive “better” or “worse”. The “similar” label is intentionally excluded, as $i-j=0$ implies identical videos.
3.2.3 Label Refinement
The label quality of the PLVD-VQA dataset inherently depends on the performance of judges. To address this dependency, we introduce an iterative label refinement framework. Specifically, once a new VQA model is trained, we treat it as an improved judge $\mathcal{J}^{\prime}$ and reapply the annotation pipeline of PLVD-VQA to iteratively refine video pair labels.
In summary, the PLVD dataset comprises $700$k annotated video pairs, partitioned into three subsets of $500$k, $100$k, and $100$k pairs, denoted as PLVD-Part1, PLVD-Part2, PLVD-Part3. The latter two subsets are dedicated to iterative label refinement via the trained model. A detailed breakdown of PLVD, including pair types and the corresponding number of videos, is provided in Table 1.
4 Proposed Method
We aim to train a VQA model on an unlabeled video dataset to compute the perceptual quality score of a video. To achieve this goal, we reformulate the VQA regression task as a classification problem that distinguishes the relative quality between pairs of videos.
4.1 Model Structure
We introduce an LMM-based VQA framework for pairwise quality ranking (LMM-PVQA). As illustrated in Fig. 3, our model comprises three components: a visual feature extractor, a text tokenizer, and an LLM decoder.
$\bm{r}$ $=\mathcal{L}(\bm{f}^{v}_{A},\bm{f}^{v}_{B},\bm{f}^{p}),$ (2)
where $\bm{r}$ is expected to belong to {“better”, “worse”} for hard ranking and {“superior”, “better”, “similar”, “worse”, “inferior”} for soft ranking.
Table 2: Performance comparison of our models against competitive opinion-unaware and supervised methods in the zero-shot setting. The best results are highlight in boldface, the second best is underlined. NA in the table indicates unavailable results. “Overall” represents the weighted average results based on the number of videos in each dataset. PLVD-P1/P2/P3 represent PLVD-Part1/ Part2/ Part3 respectively, A ✔ in the ”Label” column indicates that human-labeled data is used for training
@@ 4.1 Model Structure
Methods Training data Label SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Traning-free / Opinion-unaware VQA Methods
NIQE [38] None 0.442 0.332 0.489 0.459 0.541 0.553 0.596 0.628 0.278 0.290 0.457 0.395
IL-NIQE [72] 0.483 0.362 0.418 0.397 0.512 0.530 0.484 0.532 0.291 0.323 0.454 0.390
VIIDEO [39] 0.080 0.080 0.009 0.019 0.299 0.300 0.033 0.215 0.058 0.154 0.077 0.095
STEM [20] 0.206 0.243 0.434 0.381 0.619 0.627 0.594 0.629 0.284 0.318 0.325 0.336
TPQI [27] NA NA NA NA 0.556 0.549 0.636 0.645 0.111 0.218 0.411 0.449
BUONA-VISTA [63] NA NA NA NA 0.760 0.760 0.784 0.794 0.525 0.556 0.680 0.693
[…]
MinimalisticVQA(VII) [52] LSVQ [70] 0.861 0.859 0.740 0.784 0.843 0.841 0.757 0.813 0.775 0.779 0.817 0.830
MinimalisticVQA(IX) [52] LSVQ [70] 0.885 0.882 0.792 0.828 0.862 0.859 0.775 0.821 0.826 0.821 0.849 0.859
FAST-VQA [61] LSVQ [70] 0.880 0.880 0.781 0.813 0.859 0.854 0.826 0.845 0.730 0.747 0.838 0.849
DOVER [64] LSVQ [70] 0.878 0.866 0.782 0.813 0.874 0.869 0.817 0.840 0.771 0.781 0.842 0.845
Q-Align [65] fused [17, 11, 28, 70, 40] 0.886 0.884 0.761 0.822 0.876 0.878 0.783 0.819 0.834 0.846 0.844 0.861
[…]
Hard Ranking PLVD-P1 (500k) 0.883 0.866 0.799 0.817 0.886 0.869 0.791 0.820 0.839 0.844 0.854 0.850
Soft Ranking (Stage 1) PLVD-P1 (500k) 0.886 0.880 0.803 0.830 0.891 0.888 0.797 0.832 0.845 0.849 0.858 0.863
Soft Ranking (Stage 2) PLVD-P1/ P2 (600k) 0.887 0.880 0.802 0.830 0.893 0.892 0.798 0.833 0.856 0.859 0.859 0.864
Soft Ranking (Stage 3) PLVD-P1/ P2/ P3 (700k) 0.888 0.884 0.806 0.835 0.894 0.893 0.801 0.836 0.854 0.861 0.861 0.868

### v3

@@ 4.1.2 Ensembling Homogeneous Teachers
For video pair generation, given a pair $(x^{A},x^{B})$, each VQA model $f_{\text{weak},i}$ produces quality scores $\hat{y}^{A}_{i}$ and $\hat{y}^{B}_{i}$. We compute the mean scores $\overline{y}^{A}$ and $\overline{y}^{B}$ [^3] These weak models are trained on the same dataset and thus share the same score scale., and the score variances $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. Assuming the quality difference $\Delta=\overline{y}^{A}-\overline{y}^{B}$ follows a Gaussian distribution $\mathcal{N}(\Delta;0,\sigma^{2}_{\Delta})$ with $\sigma_{\Delta}=\sqrt{\sigma^{2}_{A}+\sigma^{2}_{B}}$, labels are assigned according to the statistical significance thresholds in (Zhu et al., 2024): “superior” if $\Delta>2\sigma_{\Delta}$, “better” if $\sigma_{\Delta}<\Delta\leq 2\sigma_{\Delta}$, “similar” if $\sigma_{\Delta}<\Delta\leq\sigma_{\Delta}$, “worse” if $2\sigma_{\Delta}<\Delta\leq-\sigma_{\Delta}$, and “inferior” if $\Delta\leq-2\sigma_{\Delta}$.
4.1.3 Integrating Heterogeneous Teachers
Another complementary approach is to diversify the teacher models in order to enrich the supervision signals. In this work, we leverage synthetic distortion simulators as specialized VQA models, which do not require human annotations for training and can be easily scaled. Concretely, we introduce three categories of synthetic distortions to emulate typical real-world degradations: spatial distortions, temporal distortions, and streaming distortions. Spatial distortions include resolution downscaling, Gaussian blur, Gaussian noise, darkening, and brightening, simulating capture-related artifacts. Temporal distortions cover jitter and stuttering, which mimic playback issues often observed in practice. Streaming distortions involve H.264 and H.265 compression, capturing compression artifacts introduced by modern media delivery platforms. The detailed simulation procedures are provided in Appendix A.3.
@@ 4.1.3 Integrating Heterogeneous Teachers
We leverage distortion severity levels (e.g., constant rate factor for compression) as pseudo-labels to infer relative quality. Given a primary video $x^{0}$ and a synthetic distortion simulator $\mathcal{S}$, we degrade $x^{0}$ across $N_{\mathcal{S}}$ severity levels to generate distorted videos $\{x_{\mathcal{S}}^{i}\}^{N_{\mathcal{S}}}_{i=1}$. Pairs $(x_{\mathcal{S}}^{i},x_{\mathcal{S}}^{j})$ are randomly sampled. Pairs with a severity difference |ij|>1|i-j|>1 are labeled as “superior” or “inferior” depending on the relative order of $i$ and $j$, while pairs with |ij|=1|i-j|=1 receive “better” or “worse”. The “similar” label is intentionally excluded, as ij=0i-j=0 implies identical videos.
Figure 6: The framework of our iterative weak-to-strong training strategy.
4.2 Iterative Weak-to-Strong Training Strategy
Within our W2S training framework, we have demonstrated that the student model can surpass its teacher models. This observation naturally motivates an iterative strategy: once a student model is trained, it can be promoted to act as a new teacher, thereby enabling another round of weak-to-strong training. Through such iterative cycles, the student progressively inherits knowledge from its predecessors while further enhancing its generalization capability. Therefore, we adopt this iterative paradigm to continually refine the student model.
From the data perspective, we expect the training samples in the next iteration to pose challenges beyond the capacity of the current teacher models, thereby further expanding the capability of the student. To this end, we introduce a difficult-sample selection strategy for both types of supervision signals in Section 4.1. Specifically, given a student model $f_{\text{w2s}}^{(i)}$ trained in the $i$-th iteration, the construction of difficult samples is straightforward for synthetic distortion pairs described in Section 4.1.2, since ground-truth labels can be directly derived from the distortion levels. We use $f_{\text{w2s}}^{(i)}$ to infer the relative quality of these pairs and select only those misclassified by the student as the training data for the $(i+1)$-th iteration.
While for the video pairs described in Section 4.1.2, no ground-truth labels are available. To address this, we adopt the group maximum differentiation (gMAD) competition framework (Ma et al., 2018) to select pairs that exhibit the largest disagreement between VQA models. Given the weak model set $\{f_{\text{weak}}^{j}\}_{j=1}^{N_{\text{weak}}}$ used to train $f_{\text{w2s}}^{(i)}$, we first partition the video pool $D_{\text{w2s}}^{(i+1)}$ into $\xi$ uniform quality levels based on the predictions of $f_{\text{weak}}^{j}$, within which videos are assumed to have similar perceptual quality. We then select pairs that are maximally differentiated by the trained student model $f_{\text{w2s}}^{(i)}$ while indistinguishable to the weak model $f_{\text{weak}}^{j}$ by
@@ 4.2 Iterative Weak-to-Strong Training Strategy
Moreover, we also reverse the roles of $f_{\text{weak}}^{j}$ and $f_{\text{w2s}}^{(i)}$ to capture cases where the student perceives similar quality but the weak model disagrees. This strategy systematically exploits the decision boundary mismatches between student and teacher models, generating informative and challenging samples that drive further improvements in next-round W2S training.
4.3 Training Strategy
We employ the standard cross-entropy loss as a baseline objective. However, weak annotations inevitably contain noise, and directly supervising the student with cross-entropy risks overfitting to erroneous labels. To mitigate this, we introduce an auxiliary confidence loss (Burns et al., 2023; Guo et al., 2024) that encourages the student to reinforce its own confident predictions, particularly when they diverge from weak labels. The overall objective is formulated as
$\mathcal{L}=(1-\lambda)\,\mathcal{L}_{\text{CE}}+\lambda\,\mathcal{L}_{\text{conf}},$ (2)
where $\mathcal{L}_{\text{CE}}$ denotes the cross-entropy loss, $\mathcal{L}_{\text{conf}}$ the confidence loss, and $\lambda$ adaptively balances label reliability against model predictions. Details of the confidence loss are provided in Appendix C.2.2.
For training data, we construct a total of 700700k annotated video pairs using the procedure described in Section 4.1.2 and Section 4.1.3. These pairs are partitioned into three subsets of 500500k, 100100k, and 100100k, denoted as $D_{\text{w2s}}^{(1)}$, $D_{\text{w2s}}^{(2)}$, and $D_{\text{w2s}}^{(3)}$, corresponding to the three stages of iterative training. A detailed breakdown of the dataset, as well as the complete training setup, is provided in Appendix A.1 and Appendix C.2.1.
Table 1: Performance comparison with state-of-the-art methods. Best and second-best results are marked in bold and underline, respectively. “Overall” represents the weighted average results based on the number of videos in each dataset.
In-domain Datasets $_{\text{test}} LSVQ$$_{\text{1080p}}$ KoNViD-1k LIVE-VQC YouTube-UGC Overall
@@ 4.3 Training Strategy
Methods SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
State-of-the-art VQA Methods - teachers
MinimalisticVQA(VII) (Sun et al., 2024) 0.861 0.859 0.740 0.784 0.843 0.841 0.757 0.813 0.775 0.779 0.817 0.830
MinimalisticVQA(IX) (Sun et al., 2024) 0.885 0.882 0.792 0.828 0.862 0.859 0.775 0.821 0.826 0.821 0.849 0.859
FAST-VQA (Wu et al., 2022) 0.880 0.880 0.781 0.813 0.859 0.854 0.826 0.845 0.730 0.747 0.838 0.849
DOVER (Wu et al., 2023a) 0.878 0.866 0.782 0.813 0.874 0.869 0.817 0.840 0.771 0.781 0.842 0.845
Q-Align (Wu et al., 2023b) 0.886 0.884 0.761 0.822 0.876 0.878 0.783 0.819 0.834 0.846 0.844 0.861
State-of-the-art VQA Methods - others
VQA2 (Jia et al., 2024) 0.878 0.872 0.794 0.821 0.881 0.880 0.785 0.830 0.811 0.823 0.847 0.854
VQAThinker (Cao et al., 2025) 0.883 0.880 0.798 0.834 0.881 0.884 0.808 0.847 0.860 0.863 0.855 0.866
Our Weak-to-Strong VQA Methods
(I): Ensembling homogeneous teachers 0.883 0.877 0.804 0.829 0.883 0.876 0.799 0.830 0.843 0.845 0.856 0.860
(II): (I) + Integrating heterogeneous teachers 0.886 0.880 0.803 0.830 0.891 0.888 0.797 0.832 0.845 0.849 0.858 0.863
(III): (II) + Confidence loss 0.885 0.881 0.803 0.831 0.890 0.891 0.797 0.833 0.849 0.856 0.857 0.865
(IV): (III) + Iterative stage W2S training 0.886 0.883 0.803 0.834 0.898 0.897 0.810 0.841 0.858 0.864 0.860 0.868
(V): (IV) + Iterative stage W2S training 0.893 0.889 0.807 0.837 0.902 0.901 0.818 0.846 0.852 0.858 0.865 0.872